OpenStudy (anonymous):

What is the sum of all the positive divisors of 72?

OpenStudy (anonymous):

$$\sigma(72)=\frac{2^4-1}{2-1}\times\frac{3^3-1}{3-1}=195$$

OpenStudy (anonymous):

In general $$\sigma(n)=\prod_{p\mid n}\frac{p^{v_p(n)+1}-1}{p-1}$$

Directrix (directrix):

If you want to go "low tech," you can get the positive divisors of 72 and add them. 1, 2, 3, 4, 6, 8, 9, 12, 18, 24, 36, 72

Directrix (directrix):

Check out the divisors online calculator at: http://www.javascripter.net/math/calculators/divisorscalculator.htm

OpenStudy (anonymous):

thanks